Circular No. 7234

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE 1999di AND 1999dj T. Puckett and A. Langoussis, Mountain Town, GA, report the dis-

covery of two apparent supernovae on unfiltered CCD frames taken with the Puckett Observatory 0.60-m automated supernova patrol telescope:

SN  1999 UT  α_{2000}  δ_{2000}  Mag.  Offset
1999di Aug. 8.38 1^{h}59^{m}54.^{s}86 +23^{o}38′22′′.7 18.0 5′′.2 E, 17′′.0 S
1999dj Aug. 8.40 1 48 10.95 +35 23 47.2 17.8 1′′.5 W, 7′′.4 S

The limiting magnitude of both discovery images was 20.0. SN 1999di in NGC 776 was confirmed on an image taken on Aug. 9.39 UT. The object does not appear on a second-generation Palomar Sky Survey (PSS II) plate (limiting magnitude 20.3) taken on 1991 Oct. 10 UT. SN 1999dj in UGC 1258 was confirmed on an image taken on Aug. 9.40 UT and does not appear on a PSS II plate (limiting magnitude 20.2) taken on 1989 Nov. 19.

COMET C/1999 P2 (SOHO) D. Lewis, University of Birmingham, reports observations of another

apparent Kreutz sugrazing comet discovered by K. Schenk in images taken with the C3 coronagraph aboard SOHO. Astrometric measurements made by Lewis (and reduced by G. V. Williams), together with orbital elements by Williams are given on MPEC 1999-P26.

1999 UT  α_{2000}  δ_{2000}
Aug. 5.362  8^{h}41.^{m}2  +15^{o}39′

COMET C/1999 N2 (LYNN) Total visual magnitudes and coma diameters: July 26.46, 7.5, - (A.

Pearce, Nedlands, W. Australia, 20 × 80 binoculars); 30.89, 7.7, 6′.0 (Jose G. de S. Aguiar, Campinas, Brazil, 11 × 80 binoculars); Aug. 1.48, 7.3, 7′ (Y. Nagai, Yamanashi, Japan, 12 × 50 binoculars); 2.07, 7.6, 3′ (J. Bortle, Stormville, NY, 20 × 80 binoculars); 2.87, 7.6, 6′ (J. Carvajal, Madrid, Spain, 0.10-m refractor); 6.90, 7.8, ∼ 4′ (R. J. Bouma, Groningen, The Netherlands, 0.16-m reflector).

CORRIGENDUM On IAUC 7233 the ‘original’ revolution period of C/1999 H1 (Lee)

should read ∼21000 years.

1999 August 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Gareth V. Williams